# Evidence of mini-jet emission in a large emission zone from a magnetically-dominated gamma-ray burst jet

### v2

@@ 1.2 Time-binned lightcurve and broad pulse fitting
| Energy range | norm | t | norm | t | |||
| (keV) | (s) | (s) | (countss) | (s) | (s) | (countss) | (s) |
@@
1.3 Profile self-similarity checking
In the Extended Data Fig 1a, we stack the time re-scaled smoothed light curves, to show that the broad features of the light curves in different energy bands share the identical shapes. The smoothed light curves are convolution between a Gaussian kernel and the original light curves. The sigma of the Gaussian kernel is 1.5 s. We use the smoothed light curve in the 500-1000 keV band as the template, and we fit the template to light curves in other bands, which are scaled by multiplying scaling factors to the their time argument . It is intuitively shown that the self-similarity of the broad profile, and this conclusion is FRED formulation independent. The scaling factor as a function of energy is plotted in the Extended Data Fig 1b, which again shows a power law dependence.
[…]
b, The scaling factor of the Gaussian smoothed light curve as a function of energy. The vertical error bars indicate the 1 uncertainties of the fitted parameters, while the horizontal error bars indicate the ranges of the energy bins.
1.4 Fast pulses
After the FRED profiles are removed from the multi-band light curves, the residuals show fast pulses (Figure 1b). It is shown clearly that the fast temporal structures are aligned in time across all energy ranges. Here, we further test this temporal alignment with cross-correlation between the residuals. In the Figure 1c, we plot the normalized correlation coefficient as a function of lag time between the residuals in a certain band and that in band 6-30 keV. As we can see, all the cross-correlation peak at lag time zero, indicating the alignment of the fast pulses.
Here we further demonstrate that some of the prominent fast structures can be identified as small pulses. For instance, we show that the spike at s can be fitted with a pulse profile. We employ our new FRED formulation to fit the pulse from 3 s to 3.75 s, and find the typical width- relation ( in the new FRED formulation) in GRB pulses. In order to show this conclusion is independent on the pulse profile formulation, we use Gaussian function as an alternative profile to fit the spike, and obtain the same width- relation ( with the Gaussian pulse formulation. See the lower panel of Extended Data Fig 2b). It can be seen in the Extended Data Fig 2a that the FRED profile results in a better fit than the Gaussian profile. The fitted parameters are list in Extended data Table 2. It is also interesting to observe that this small pulse does not show spectrum lag of its peak time as the broad pulse. We demonstrate this by cross-correlating between the small pulses in multi-bands and that in 30-70 keV channel. The lag time together with their uncertainties as a function of energy is shown in the upper panel of Extended Data Fig 2b, where all lag times are almost zero, and we can clearly conclude that the lag time does not depend on energy like the broad pulse.
@@ 1.4 Fast pulses
| Energy range | norm | norm | t | |||
| (keV) | (s) | (s) | (countss) | (s) | (countss) | (s) |
@@
[…]
The dip at s is another noticeable feature of the multi-band light curves. A natural attempt is to attribute such a dip to some sort of absorption or geometrical blocking. We define the effective “optical depth" in different energy bands as:
where is the net count rate at the bottom of the dip, which is found with a negative Gaussian fitting superposed on the FRED of broad pulse fitting from 17 s to 19.5 s (Extended Data Fig 3a); and is the net count rate of the broad feature. The results of the fitting with a negative Gaussian is tabulate in Extended data Table 3. It is intriguing to find that there is a power law energy dependence of the effective optical depth, and the best fit power law index is (Extended Data Fig 3c). This energy dependence of the optical depth challenges the absorption picture, for there is no known absorption mechanism whose cross section proportional to energy to the order of with a break at about 300 keV.
[…]
Here we view the dip as gaps between two successive fast temporal structures. Therefore, we fit a pulse profile (new FRED formulation) and a rising edge of another pulse (Norris05) to the dip at each energy bands from 17 s to 19.5 s (Extended Data Fig 3b). The fitted parameters are list in Extended data Table 4. The time scale of the earlier pulses as a function of energy is plotted in Extended Data Fig 3d, which shows clearly a typical power law energy dependence of pulse width.
@@
| pluse before gap | pluse after gap | additional constant | |||||
| Energy range | t | norm | t | norm | constant | ||
| (keV) | (s) | (s) | (countss) | (s) | (s) | (countss) | (countss) |
@@
1.5 Physical implications
Here we demonstrate that the observed features presented in this work cannot be explained within the framework of the internal shock (IS) model, but can be naturally reproduced with models involving many mini-jets in a single expanding magnetized fluid, as envisaged in the ICMART model.
In the standard IS picture [5], a random distribution of the shell bulk Lorentz factor results in internal shocks at random but similar distances from the central engine. These collisions generate a series of independent fast pulses. In this case, any long term overall trend of the light curve would have to reflect the activity history of the central engine. The same time history would be reflected in different energy bands. Even if each short pulse may have a energy-dependent behavior in the timescale of the short pulse itself, there is no reason to show a global behavior of shape-energy dependence (see the cartoon in Extended Data Fig. 4 for an illustration). To further demonstrate this, we conducted a series of simulations within the framework of the standard IS model with the following setting: We simulated the ejection of 50 shells from the central engine, with their initial thicknesses, Lorentz factors and masses sampled from random distributions in the logarithmic space. The Lorentz factors vary from 100 to 1000, the mass ranges from to , and the initial thickness ranges from to , all in cgs units. The central engine’s ejection times are randomly selected from a linear distribution, with values ranging from to seconds in the rest frame of the GRB central engine. Upon each collision, two shells merge, and the merged shell parameters (, , ) are adopted as the new values for the remaining shell. To ensure tracking future collisions, the “effective" ejection time of this new shell is redefined as , where and is the Lorentz factor of the merged shell. The simulation then restarts with one less shell. This process is repeated for each collision, allowing the code to monitor all collision/merging events for any arbitrarily designed central engine activity [6, 39]. The results of the simulation is presented in Extended Data Fig. 7a, which confirm our expectation that there is no shape-energy dependence of the overall lightcurve. In the IS model, the energy-dependent lag and width generally apply to each pulse. Even if there is a stretching factor in low energies to make pulses broader, the stretching factor scales with the duration of short pulses. There is no way to reproduce a large stretching factor in the scale of the broad pulse by superposing these short pulses.
In an variant IS model[7, 40, 41], a separate shock at a larger distance () is generated to make a slow pulse component that overlaps with those fast pulses. The overall lightcurve is dominated by the slow pulse and shows the shape-energy dependence (as illustrated in Extended Data Fig. 5). However, in this case, there should be an underlying slow pulse emission component, which is in direct contradiction with the observations presented in this work. The corresponding simulation results are presented in Extended Data Fig. 7b, which does not reproduce the observations, especially the deep dip at s. More details of the above-mentioned two IS simulations will be reported in a separated paper.
Within the framework of the ICMART model, collisions between a pair of highly-magnetized shells with different Lorentz factors would distort ordered magnetic field and trigger fast magnetic seeds, which will induce relativistic magnetohydrodynamic (MHD) turbulence. Numerical simulations have demonstrated that MHD turbulence has the ability to generate random local magnetic reconnection events [42, 43, 44]. These reconnection events would convert magnetic energy in the fluid to the particle energy, which would further distort magnetic fields, resulting in a reconnection cascade. The particle energy would then be transfer to the kinetic energy of the bulk through adiabatic cooling and to the observed photon energy through synchrotron radiation. The overall trend of the light curve corresponds to the growth of the number of magnetic reconnection events and a global curvature effect due to high-latitude emission after the line-of-sight emission ceases[15, 16]. Each mini-jet emission contributes to a fast pulse and their random latitudes and orientations cause the high variability of the lightcurve (see illustration in Extended Data Fig. 6). Dips could occur at random instances when the waiting time of pulses from successive mini-jets is long enough.
As the fluid expands, the gloabl magnetic field strength in the emission region naturally decreases. This leads to a global decline of synchrotron peak energy and causes a well-defined shape-energy dependence behavior, including spectral lags and global hard-to-soft evolution[17, 18]. Furthermore, photons from reconnection events at high latitudes arriving at later times after the line-of-sight emission is over lead to a redder spectrum because of its smaller bulk Doppler factor[16]. All these factors ensure an overall trend of a softer-wider/softer-later behavior.
@@ 1.5 Physical implications
We have presented a series of simulations to test this picture. We assume a pair of magnetized shells collide and merge as one fluid after a collision at cm. The mass of the fluid is assumed to be g, the initial Lorentz factor is the initial magnetization factor is (i.e. the comoving-frame magnetic energy is times of the rest-mass energy of the fluid). The number of reconnection events increases by a factor of during the emission period and the total number of the reconnection events within the cone is assumed to be . The relatively small number of reconnection events was chosen to ensure the occurrences of dips in most of the realizations. Each reconnection event has a size of cm, a random latitude within and a random oriention within . In Extended Data Fig. 7 c, we show one realization of the simulations that the following features consistent with the observations: 1) the lightcurve can be decomposed of many short pulses related to mini-jet emission; 2) the overall shape of the lightcurve resembles a single pulse, but can possess a very deep dip; 3) the overall single slow pulse has a significant shape evolution across energy bands. This simulation demonstrates that the ICMART model can well reproduce the observed behavior of GRB 230307A. Detailed simulations will be presented in a separate paper.
[…]
b, The results for the modified internal shock model that invokes superposition a large-radius internal shock and many small-radii internal shocks. The overall light curve is dominated by the emission from the large-radius internal shock, with no hope to produce dips in the energy-dependent lightcurves;
c, The results for the ICMART model. Both aligned short pulses and a global shape-energy dependence can be repduced. The shaded area indicates the location of a reproduced lightcurve dip due to the lack of mini-jet emission at that time.
Data Availability
The processed data are presented in the tables and figures of the paper, which are available upon reasonable request.
Code Availability
Upon reasonable requests, the code (mostly in Python) used to produce the results and figures will be provided.

### v4

@@ 3.2 The self-similarity of the overall profile
| Energy range | norm | ts | norm | ts | |||
| (keV) | (s) | (s) | (countss-1) | (s) | (s) | (countss-1) | (s) |
@@
The “self-similarity” inspired us to propose another formulation of FRED profile, with one less parameter than that of Norris05:
We fit the multi-band light curves again with the new FRED formulation. The results of fitting are listed in Table 1. The sole time scale parameter shows a similar energy dependence on energy, , from 6 to 300 keV, and shows a hint of shallower energy dependence above keV (Figure 2). In the new formulation, one can define peak time as , where we denote the in the channel with . The width of the FRED profile also scales as . Therefore, the found dependence can naturally result in the and relations.
3.3 The fast components
We subtract the slow trend described by equation 2 with the best-fit parameters, to study the fast varying structures. From the residuals of the light curves (Figure 5), one can see that there are many rapidly-varying short pulses. One noticeable feature is that even the broad pulse has a clear energy dependence, the short time spikes and dips appear to align at the same instances across different energy bands.
We further demonstrate this alignment of the fast temporal features by cross-correlations among multi-band light curve residuals. As shown in the Figure 6, the peaks are perfectly aligned across the full energy band. This immediately excludes the attempt to explain the self-similar stretchable feature of the multi-band light curves as relativistic time-dilation, as a time-dilation effect would stretch both slow and fast varying pulses together.
Here we further demonstrate that some of the prominent fast structures can be identified as small pulses without an underlying slow component. For instance, we show that the spike at s can be fitted with a pulse profile. We employ our new FRED formulation to fit the pulse from 3 s to 3.75 s, and find the typical width- relation ( in the new FRED formulation) in GRB pulses. In order to show this conclusion is independent on the pulse profile formulation, we use Gaussian function as an alternative profile to fit the spike, and obtain the same width- relation ( with the Gaussian pulse formulation. See the lower panel of Figure 8. It can be seen in the Figure 7 that the FRED profile can also be used to fit the individual pulse, with better representation of its asymmetry. The fitted parameters are list in Table 2. It is also interesting to observe that this small pulse does not show spectrum lag of its peak time as the broad pulse. We demonstrate this by cross-correlating between the small pulses in multi-bands and that in 30-70 keV channel. The lag time together with their uncertainties as a function of energy is shown in the upper panel of Figure 8, where all lag times are almost zero, and we can clearly conclude that the lag time does not depend on energy like the broad pulse.
@@ 3.3 The fast components
| Energy range | norm | norm | ts | |||
| (keV) | (s) | (s) | (countss-1) | (s) | (countss-1) | (s) |
@@
3.4 The dip at around 18s
The dip at s is another noticeable feature of the multi-band light curves. A natural attempt is to attribute such a dip to some sort of absorption or geometrical blocking. We define the effective “optical depth” in different energy bands as:
where is the net count rate at the bottom of the dip, which is found with a negative Gaussian fitting superposed on the FRED of broad pulse fitting from 17 s to 19.5 s (Figure 9); and is the net count rate of the broad feature. The results of the fitting with a negative Gaussian is tabulate in Table 3. It is intriguing to find that there is a power law energy dependence of the effective optical depth, and the best fit power law index is (Figure 10). This energy dependence of the optical depth challenges the absorption picture, for there is no known absorption mechanism whose cross section proportional to energy to the order of with a break at about 300 keV. Beside, if we were to explain the dip as absorption, its depth and transience require that the absorbing object has a angular size comparable to that of the emission region, and a very high transverse angular velocity. That will either result in a superluminal transverse velocity of the blocking object, or highly coincidental absorption in our local.
@@ 3.4 The dip at around 18s
A reasonable explanation of the dip is the gap between two successive fast temporal structures. In order to demonstrate, we fit a pulse profile (new FRED formulation) and a rising edge of another pulse (Norris05) to the dip at each energy bands from 17 s to 19.5 s (Figure 11). The fitted parameters are list in Table 4. The time scale of the earlier pulses as a function of energy is plotted in Figure 12, which shows clearly a typical power law energy dependence of pulse width.
| pulse before gap | pulse after gap | additional constant | |||||
| Energy range | ts | norm | ts | norm | constant | ||
| (keV) | (s) | (s) | (countss-1) | (s) | (s) | (countss-1) | (countss-1) |
@@
3.5 Summary of the observation features
Here we summarize the intriguing features from observation: (a) the overall light curve resembles a classic “single pulse”, which is wider and peaks later in lower energies in a self-similar way; and (b) there is no underlying elementary single slow pulse and the apparent broad profile is composed with many fast pulses.
In order to better demonstrate that the observed profile-energy dependence cannot be attributed to an underlying broad pulse, we perform the following analysis: we downscale the best-fit FRED-profiles in each energy bands, so that there are no negative residuals in the high time-resolution light curves. The rescaled broad FRED are the maximum allowed underlying broad pulses in each energy bands, we plot it in the Figure 13. The found rescaling factor is the fraction of fluence attributed from this broad pulse. We can see that the rescaling factor in most energy bands are found to be neglegible, suggesting that the broad pulse-energy dependence cannot be attributed to an underlying broad pulse.
In other words, the individual pulses work in concert with each other to make the overall profile-energy dependence.
4 Physical implications
The above features found in this burst indicate that individual fast pulses are correlated in time and energy. The implication of that is, the emission sites of individual pulses should be connected casually. It provides significant insights into the prompt emission mechanism and the composition of the jet.
4.1 The ICMART model
In the ICMART model (Zhang & Yan, 2011; Zhang & Zhang, 2014; Shao & Gao, 2022), individual fast pulses are emitted from mini-jets launched by local magnetic reconnections. All mini-jets are in the same expanding magnetically-dominated fluid (see illustration in Figure 14). The rise and fall of the FRED-like profile is mostly dictated by the rate of growth of the number of reconnection events (Yi et al., 2024). During the evolution of the broad pulse, the magnetic field strength decays as the whole fluid expands (Uhm & Zhang, 2014). It leads to rolling down of the characteristic synchrotron emission frequency with time (Uhm & Zhang, 2016; Uhm et al., 2018). The soften of the spectra over time leads to the observed profile-energy dependence 333See Sun et al. (2024) for the spectrum time evolution analysis of this burst.. With certain spectrum evolution, a self-similar profile may be reproduced (Yi et al., 2024). Within this picture, the observed prompt emission originates from a single episode of central engine activity, which is consistent with the merger origin of GRB 230307A (Gillanders et al., 2023; Sun et al., 2024; Yang et al., 2024).
@@ 4.1 The ICMART model
In order to better illustrate the idea, we conduct ICMART simulations and present the results. The general settings of the simulation are as follows: We assume a pair of magnetized shells collide and merge as one fluid after a collision at cm. The mass of the fluid is assumed to be g, the initial Lorentz factor is the initial magnetization factor is (i.e. the comoving-frame magnetic energy is times of the rest-mass energy of the fluid). The number of reconnection events increases by a factor of during the emission period and the total number of the reconnection events within the cone is assumed to be . The relatively small number of reconnection events was chosen to ensure the occurrences of dips in most of the realizations. Each reconnection event has a size of cm, a random latitude within and a random orientation within . For each reconnection, the electrons are assumed to be accelerated to power law distribution with index -2.8. The electrons cool down and radiate via synchrotron radiation. We calculate the cooling and radiation process according to Uhm & Zhang (2014), and superpose the resulting synchrotron spectrum of each mini-jet to obtain the observed spectrum. In Figure 15, we show the light curves reproduced by the simulation that is consistent with the key observation features: 1) the light curve can be decomposed of many short pulses related to mini-jet emission; 2) the overall shape of the light curve resembles a single pulse, but can possess a very deep dip; 3) the overall single slow pulse has a significant shape evolution across energy bands. This simulation demonstrates that the ICMART model can well reproduce the observed behaviour of GRB 230307A. Detailed simulations will be presented in a separate paper.
4.2 Standard internal shock model
The standard internal shock model, on the other hand, attributes short-time pulses as a consequence of collisions of many pairs of shells (see Figure 16 for an illustration). Within such a scenario, the overall trend is determined by the history of the central engine activity. Light curves in different bands should share this same history and should not display a global energy-dependent behaviour as observed.
To better demonstrate this, we conducted simulations within the framework of the standard IS model with the following settings: We simulated the ejection of 50 shells from the central engine, with their initial thicknesses, Lorentz factors and masses sampled from random distributions in the logarithmic space. The Lorentz factors vary from 100 to 1000, the mass ranges from to , and the initial thickness ranges from to , all in cgs units. The central engine’s ejection times are randomly selected from a linear distribution, with values ranging from to seconds in the rest frame of the GRB central engine. Upon each collision, two shells merge, and the merged shell parameters (, , ) are adopted as the new values for the remaining shell. To ensure tracking future collisions, the “effective” ejection time of this new shell is redefined as , where and is the Lorentz factor of the merged shell. The simulation then restarts with one less shell. This process is repeated for each collision, allowing the code to monitor all collision/merging events for any arbitrarily designed central engine activity (Kobayashi et al., 1997; Maxham & Zhang, 2009). The spectra for each pulse are simulated empirically (see details in a separate paper (Moradi et al., 2024)). The results of the simulation is presented in Figure 17, where there is no shape-energy dependence of the overall light curve as we expected. In the IS model, the energy-dependent lag and width generally apply to each pulse. Even if there is a stretching factor in low energies to make pulses broader, the stretching factor scales with the duration of short pulses. There is no way to reproduce a large stretching factor in the scale of the overall profile by superposing these short pulses.
4.3 Modified internal shock model with another internal shock at large radius
A modified internal shock model (see Figure 18) may invoke another internal shock at a much larger emission radius (Daigne & Mochkovitch, 1998b, 2003; Bošnjak & Daigne, 2014) defined by , where s is the timescale of the broad pulse. However, within such a picture, the observed emission should be the superposition between the fast and slow components, which requires that fast pulses only grow on top of a dominating the slow component. It is not supported by the existence of the very deep, achromatic dip at around 18 s.
We also conduct a simulation under this variant IS model: a separate shock at a larger distance () is generated to make a slow pulse component that overlaps with those fast pulses. We consider synchrotron radiation process in IS model to determine the spectrum evolution of the long pulse, using the analytical solution proposed by Daigne & Mochkovitch (2003), involving a relativistic wind where a slower shell () decelerates a faster portion of the flow () characterized by a constant rest-frame mass flux. The overall light curve is dominated by the slow pulse and shows the shape-energy dependence (as illustrated in Figure 18). However, in this case, there should be an dominating underlying slow pulse emission component, which is in direct contradiction with the observations presented in this work. The corresponding simulation results are presented in Figure 19, which does not reproduce the observations, especially the deep dip.
In summary, any other attempt trying to model the prompt emission of this bursts within such frameworks should manage to address all the aforementioned challenges.
5 Summary
In this paper, we demonstrate that the light curve of GRB 230307A can be decomposed into several individual pulses, while the overall behavior of these individual pulses (slow varying trend) resembles a self-similar broad single pulse. While the rapid variability shows alignment in all energy bands, the broad pulse has a well-defined energy-dependence manner. In order to reach the above observational conclusion solidly, we conduct the following analysis:
We fit parameterized FRED profiles to the overall light curves in different energy bands, and found that the width and peak time of the profile decrease with the energy in a power law. This dependence was found to saturate at keV;
We found that the smoothed overall shape of the light curve in each band is essentially a time-stretched copy of that in the other bands;
After subtraction of the over all trend of the light curve, we show that the residual variations superimposed on the best-fit ”self-similar” temporal profile do not undergo the same temporal stretching, but are instead fixed in time regardless of the band.
We found that some prominent fast components can be fitted with individual pulses, such as the pulse at around 3.5 s, and the dip at around 18 s, without the need of an underlying slow components;
The deep dip at around 18 s significantly constrained any possible elementary slow components, to which the observed overall profile-energy dependence can be attributed.
The above mentioned features indicate that the prompt emission of this burst is from many individual emitting units that are casually linked in a emission site at a large distance from the central engine. We elaborated and demonstrated with simulations that, this scenario is consistent with the picture of many mini-jets due to local magnetic reconnection events in a large emission zone far from the GRB central engine (as envisaged in the ICMART model) but raises a great challenge to the internal shock-like models that attribute all variability components to collisions among different shells.